# COMPASS: Robust Feature Conformal Prediction for Medical Segmentation Metrics

### v1

@@ 2.3 COMPASS Calibration and Inference
Regardless of the features used to compute direction $\Delta_{i}$, the calibration and inference procedures for COMPASS follow the standard CP protocol. On the calibration set $\{(x_{i},y_{i})\}_{i\in D_{\mathrm{cal}}}$, our goal is to find the non-conformity score $R_{i}$ for each sample. We first define the prediction interval for a given perturbation magnitude $\beta\geq 0$ as:
$S_{\beta}(x_{i}):=[h(g(\hat{z}_{i}-\beta\Delta_{i})),h(g(\hat{z}_{i}+\beta\Delta_{i}))].$
The score $R_{i}$ is then the minimal magnitude that ensures coverage of the true value $y_{i}$:
$R_{i}:=\inf\{\beta\geq 0:y_{i}\in S_{\beta}(x_{i})\}.$
To find this score efficiently, we perform a binary search with $\mathcal{O}(\log N)$ complexity. From the set of calibration scores $\{R_{i}\}_{i=1}^{n}$, we then compute the calibrated quantile $\hat{\beta}$ with the standard finite-sample correction:
$\hat{\beta}:=Q\left(\{R_{i}\}_{i=1}^{n},\frac{\lceil(1-\alpha)(n+1)\rceil}{n}\right).$
At inference time, for a new sample $x_{n+1}$, we compute its direction $\Delta_{n+1}$ and form the final prediction interval by applying the calibrated quantile $\hat{\beta}$:
$S_{\hat{\beta}}(x_{n+1}):=[h(g(\hat{z}_{n+1}-\hat{\beta}\Delta_{n+1})),h(g(\hat{z}_{n+1}+\hat{\beta}\Delta_{n+1}))].$
By Theorem 1, this interval is guaranteed to achieve the desired marginal coverage. For a full outline of the algorithm, including the asymmetric case, refer to Appendix C.
While symmetric perturbations are effective when the metric responds similarly to positive and negative perturbations, we often observe an asymmetric relationship when the metric responds differently in positive and negative directions. In such cases, an asymmetric calibration is necessary to construct a more adaptive and efficient interval. We achieve this by finding separate non-conformity scores for the upper and lower bounds. We discuss this asymmetric version of COMPASS and provide an equivalent algorithm in Appendix B.
@@ 2.4 Weighted COMPASS for Distribution Shifts
To address this limitation, we employ Weighted Conformal Prediction (WCP) [22, 23] to restore the coverage guarantee by re-weighting the calibration samples. The weight for a calibration sample $X_{i}$ is ideally the density ratio $w(X_{i})=p_{\mathrm{test}}(X_{i})/p_{\mathrm{cal}}(X_{i})$. In practice, this ratio is unknown and typically estimated by training an auxiliary classifier $\mathcal{A}:\mathcal{X}\to[0,1]$ to distinguish between samples from the calibration and test sets. The predicted probability $\hat{p}(x)=\mathcal{A}(x)$ that a sample $x$ belongs to the test set is then used to compute the weights, effectively adjusting the calibration procedure to account for the distribution shift. We now extend Theorem 1 to the weighted setting, providing a formal coverage guarantee under distribution shift. The setup remains identical.
@@
This is a direct application of WCP [22]. The validity of this framework applies to any valid non-conformity score. Our scores are valid as they are from a deterministic function of the samples and the pre-trained model. When the weights represent the true density ratio, the weighted empirical distribution of the calibration scores is an unbiased estimator of the test score distribution. Thus, the coverage guarantee from the original theorem directly applies. ∎
In practice, the true oracle is not available, and the coverage guarantee holds approximately, with the quality of the approximation depending on the accuracy of the density ratio estimates. To correct for the induced distribution shift, we explore three correction strategies: 1) Class: available ground truth class labels as features and weights are computed directly from the known class prevalences, 2) Latent: Model’s latent representations summed on the spatial dimension as features [24], and 3) Jacobian: model’s internal geometric sensitivity (jacobians) summed as features.
@@ 3 Experiments
We evaluated COMPASS across four medical image segmentation tasks: 1) segmentation on H&E histopathology images from the EBHI dataset (H&E) [18], 2) skin lesion segmentation on dermoscopic images from the HAM10000 dataset (Skin Lesion) [19], 3) thyroid nodule segmentation from the TN3K dataset (Nodule) [20], and 4) gastrointestinal polyp segmentation (PolyP) [21]. We trained all models using the standard U-Net architecture from MONAI [25, 26]. We focused on segmented object size (area) as the downstream clinical metric of interest. Object size is among the most widely adopted quantitative biomarkers across diverse clinical applications [27, 28]. We repeated each experiment over 100 randomized splits to compute average coverage and interval sizes. We provide details on architecture, preprocessing, training, calibration, and testing details in Appendix E. We considered three types of calibration strategies:
1. Output-space calibration: These include split conformal prediction (SCP) [29], Conformalized Quantile Regression (Output-CQR) [30], and locally adaptive conformal prediction (Local CP) [29, 31, 32].
2. End-to-end calibration: We use Conformalized Quantile Regression (E2E-CQR) [24] on a model trained to directly produce pixel-wise lower and upper bounds using the Tversky loss [33].
3. Feature-space calibration: These include COMPASS-L, which directly calibrates the model’s final pre-activation outputs (logits), and COMPASS-J, which constructs prediction intervals by perturbing latent features along dominant directions. We omitted the original FCP method [15] from our main empirical comparison because we found its core adversarial search procedure to find non-conformity scores to be computationally intractable for large feature spaces and fails to reliably converge. We instead offer a comparison against a conceptual oracle benchmark for FCP in Appendix D: because FCP prediction set is defined by an $L_{p}$ ball in the latent space [15] with bounds computed using Linear Relaxation based Perturbation Analysis (LiRPA) [34], we empirically find the minimal radius of the $L_{p}$ ball, which, when propagated through the decoding function using LiRPA, achieves the target coverage for the final metric.
Table 1: COMPASS achieves more efficient interval sizes compared to baseline methods across different target coverages. For 4 datasets and 100 random splits, we compare interval lengths at target coverages at $\alpha=\{0.15,0.1,0.05\}$. We show the output space, end-to-end, and feature calibration methods in red, blue, and green. The shortest mean interval lengths are bolded. For empirical coverages, see Appendix D.
Interval Size (Mean±STD, Pixels2)
@@
H&E 0.05 4635±569 4542±449 5161±465 6211±598 5392±338 5805±732
0.1 2512±160 2655±205 2929±312 3764±322 3637±237 3467±265
0.15 2364±145 2513±145 2652±169 3182±249 3082±284 2733±199
Skin Lesion 0.05 1590±61 1633±84 2570±211 3953±204 10936±55 3296±236
0.1 1149±41 1170±47 1344±80 2580±166 4911±41 1828±138
0.15 920±24 932±26 935±40 1866±62 2470±36 1118±77
Nodules 0.05 3184±190 3293±208 4075±320 3810±234 7560±49 4616±375
0.1 2388±130 2469±141 2810±184 3223±99 5450±48 3095±221
0.15 1940±106 2005±138 2197±181 2706±148 4066±88 2381±154
PolyP 0.05 5061±497 5358±563 8934±843 12071±2197 7618±820 8305±574
0.1 3940±262 4101±286 6329±584 6324±926 4887±459 6194±550
0.15 3228±191 3414±254 4593±496 4391±410 3951±295 4533±405
@@ 3.1 Standard Conformal Prediction
First, we qualitatively find that COMPASS finds semantically meaningful directions in the latent space that correspond to monotonic changes in metric (Figure 2) for all applications. Next, we compare interval size across methods and datasets for $\alpha=\{0.15,0.1,0.15\}$ (Table 1). We find that output space methods (SCP, CQR, Local) generally achieve valid coverage (Table 2 in Appendix D) but yield longer intervals. End-to-end CQR provides tighter intervals. COMPASS (-J and -L) achieves valid coverage with a significantly reduced interval size compared to E2E-CQR. Moreover, we observe an efficiency benefit of using deeper representations (COMPASS-J) compared to logits (COMPASS-L). COMPASS-J generally achieves the shortest interval lengths overall. This highlights the efficiency of COMPASS methods.
Figure 2: As latent features are shifted along the pertubation directions $\Delta$, the induced segmentation volumes monotonically expand. Using COMPASS-J and for a sample from each dataset, we show a grid of segmentation contours (in red) with perturbation magnitudes $\beta$ targeted at -20%, -10%, 0% (original prediction), +10%, +20% change in area ($\delta A)$. We provide more examples in Appendix G.
@@ 3.2 Weighted Conformal Prediction
To empirically validate the effectiveness of WCP and compare different weighting strategies, we design an experiment with a controlled, adversarial label shift for datasets that include class labels, namely H&E and Skin Lesion. Using the same architecture and training setup as the unweighted case, we create two scenarios where unweighted methods are expected to fail, thereby highlighting the ability of WCP to restore the target coverage and efficient interval lengths. The training set is the same as the unweighted case. The calibration set is constructed to be 1) “easy” by undersampling a difficult class and 2) “hard” by undersampling the easy class. The test set is constructed to be 1) “hard” by oversampling a difficult class and 2) easy by undersampling the difficult class. For H&E, we place 60% and 40% of the Adenocarcinoma (easy) samples into the calibration and test sets. For Skin Lesion, we place 70% and 30% of the melanocytic nevi (hard) samples into the calibration and test sets. See Appendix E for the calibration and testing class distributions.
For output-space calibration methods, we used weights based on the ground truth class labels, which we consider as approximately “oracle” weights. Note that the theoretically perfect oracle weights are defined by the density ratio, which precisely corrects for the change in the full, high-dimensional distribution of features between the calibration and test sets. For end-to-end and feature space calibration, we used the latent features [24, 35, 36] and Jacobians to train auxiliary classifiers using gradient boosting machines [37] to distinguish between calibration and test sets. For end-to-end, we do not use Jacobian weights consistent with prior work [24].
We plot the distribution of interval size and mean coverage for unweighted, oracle, feature, and Jacobian weighted methods with $\alpha=0.1$ (Figure 3). As expected, unweighted methods do not match the target coverage. The coverages for H&E and Skin Lesion tend to be under or overcovered, which is expected from the adversarial split procedure. We find that using class, feature, and Jacobian weightings allows target coverage recovery. In each weighting method that achieves target coverage, COMPASS (specifically COMPASS-J) achieves the shortest interval lengths. The slight undercoverage for feature-weighted E2E-CQR for H&E and class weighted methods in Skin Lesion can be attributed to the unavailability of the true oracle weights. In spite of this, COMPASS methods are the closest to target coverage. These results highlight the strength of COMPASS for correcting for covariate shifts. We show a table of results in Appendix D.
@@ 4 Discussion
We introduced COMPASS, a unified framework to generate efficient, metric-based prediction intervals by leveraging the inductive biases of neural networks. Under the assumption of nestedness and exchangeability, COMPASS creates practical feature-space calibration by perturbing a model’s intermediate representations along low-dimensional subspaces maximally sensitive to the target metric. Across four medical segmentation tasks for area estimation, COMPASS achieves valid coverage while producing significantly tighter intervals than traditional output-space and end-to-end baselines for both standard and weighted CP. We discuss several further points below.
COMPASS has several attractive properties. First, it produces instance-adaptive intervals. For COMPASS-J, the perturbation direction for any new sample is found by projecting its Jacobian onto the principal component component and reconstructing the direction. Second, COMPASS performs better than naive output-space calibration. In tasks like segmentation, the final metric is a complex non-linear function of the model’s output (the pixel mask). Simply adding a margin to a final object size prediction is a crude approximation that fails to capture how the object size is actually derived from the underlying segmentation. In contrast, COMPASS exploits the model’s internal structure and spatial understanding to directly manipulate and calibrate the metric. COMPASS has more information and degrees of freedom, which allows it to achieve more efficient interval sizes. Third, COMPASS performs better than methods trained using pixel-level losses such as E2E-CQR. Such approaches optimize a proxy (pixel quantiles) for the downstream metric, that may or may not translate into accurate intervals for the final metric. On the other hand, COMPASS is directly calibrated to the final metric, and is therefore more applicable to practical clinical use cases.
Finally, COMPASS is naturally applicable to various segmentation architectures. We also present the interval lengths and coverage for SegResNet [38] in Appendix D. COMPASS produces the most efficient interval lengths on the majority of datasets across $\alpha=\{0.05,0.1,0.15\}$ for SegResNet.
@@
COMPASS’s performance is fundamentally dependent on the quality of the pre-trained model’s representations. Our framework presupposes that the model has learned an inductive bias where the chosen feature space has a reasonably monotonic and sensitive control over the downstream metric. If the features are poorly aligned with the metric, the resulting perturbation direction may be inefficient or non-monotonic, leading to wide or invalid intervals. Furthermore, as mentioned in prior work [24], WCP with feature or jacobian weights, is effective for moderate distribution shifts where the calibration and test distributions have significant overlap in the feature space. However, for large shifts, the estimated weights may be inaccurate, which can compromise the validity of the final interval.
5 Conclusion
We introduce COMPASS, a unified framework to generate efficient, metric-based prediction intervals by leveraging the inductive biases of neural networks. Under the assumption of nestedness, COMPASS creates practical feature-space calibration by perturbing a model’s intermediate representations along low-dimensional subspaces maximally sensitive to the target metric. Across four medical segmentation tasks for area estimation, COMPASS achieves valid coverage while producing significantly tighter intervals than traditional output-space and end-to-end baselines for both standard and weighted CP.

### v2

@@ 2.3 COMPASS Calibration and Inference
The goal of COMPASS is to find the smallest symmetric perturbation magnitude $\beta$ such that the entire interval $S_{\beta}(x)$ contains the ground truth value $y$. In general, computing $S_{\beta}(x)$ requires finding the minimum and maximum of the perturbed metric response $m_{x_{i}}(b),b\in[-\beta,\beta]$. When the metric response is non-monotonic, computing the extrema requires a full discretized sweep across the perturbation range. Performing this sweep at every step during calibration makes the envelope computation expensive because each candidate $\beta$ requires many forward passes to map out the complete response curve for each sample.
However, when $m_{x_{i}}(b)$ is monotonic in $b$, the conservative envelope collapses to evaluations at the endpoints: $S_{\beta}(x_{i})=\big[m_{x_{i}}(-\beta),\;m_{x_{i}}(+\beta)\big]$. Thus, the interval can be computed with only two forward passes, giving a practical and efficient implementation of COMPASS. A perturbation sweep is still required, but only once per sample, solely to verify monotonicity rather than to repeatedly compute the envelope during calibration. As we will see in Figures 2 and 12, this monotonicity condition holds across all of our experiments.
Once $S_{\beta}(x_{i})$ can be evaluated, either via the conservative full sweep or the endpoint method, the non-conformity scores $\{R_{i}\}_{i=1}^{n}$ follow directly from Equation 3. The conformalized quantile is:
$\hat{\beta}=Q\!\left(\{R_{i}\}_{i=1}^{n},\frac{\lceil(1-\alpha)(n+1)\rceil}{n}\right),$
and at test time we return the interval $S_{\hat{\beta}}(x_{n+1})$.
For a full outline of the algorithms, refer to Appendix C. Furthermore, while symmetric perturbations are effective when the metric responds similarly to positive and negative perturbations, we often observe an asymmetric relationship when the metric responds differently in positive and negative directions. In such cases, an asymmetric calibration, where we find separate non-conformity scores for the upper and lower bounds, is necessary to construct a more adaptive and efficient interval. We discuss this asymmetric version of COMPASS and provide an equivalent algorithm in Appendix B.
@@ 2.4 Weighted COMPASS for Distribution Shifts
To address this limitation, we employ Weighted Conformal Prediction (WCP) [Tibshirani et al., 2019, Barber et al., 2023] to restore the coverage guarantee by re-weighting the calibration samples. The weight for a calibration sample $X_{i}$ is ideally the density ratio $w(X_{i})=p_{\mathrm{test}}(X_{i})/p_{\mathrm{cal}}(X_{i})$. In practice, this ratio is unknown and typically estimated by training an auxiliary classifier $\mathcal{A}:\mathcal{X}\to[0,1]$ to distinguish between samples from the calibration and test sets. The predicted probability $\hat{p}(x)=\mathcal{A}(x)$ that a sample $x$ belongs to the test set is then used to compute the weights, effectively adjusting the calibration procedure to account for the distribution shift. We now extend Theorem 1 to the weighted setting, providing a formal coverage guarantee under distribution shift. The setup remains identical.
@@
This is a direct application of WCP [Tibshirani et al., 2019]. The validity of this framework applies to any valid non-conformity score. Our scores are valid as they are from a deterministic function of the samples and the pre-trained model. When the weights represent the true density ratio, the weighted empirical distribution of the calibration scores is an unbiased estimator of the test score distribution. Thus, the coverage guarantee from the original theorem directly applies. ∎
In practice, the true oracle is not available, and the coverage guarantee holds approximately, with the quality of the approximation depending on the accuracy of the density ratio estimates. To correct for the induced distribution shift, we explore three correction strategies: 1) Class: available ground truth class labels as features and weights are computed directly from the known class prevalences, 2) Latent: Model’s latent representations summed on the spatial dimension as features [Lambert et al., 2024], and 3) Jacobian: model’s internal geometric sensitivity (jacobians) summed as features.
@@ 3 Experiments
We evaluated COMPASS across four medical image segmentation tasks: 1) segmentation on H&E histopathology images from the EBHI dataset (H&E) [Hu et al., 2023], 2) skin lesion segmentation on dermoscopic images from the HAM10000 dataset (Skin Lesion) [Tschandl et al., 2018], 3) thyroid nodule segmentation from the TN3K dataset (Nodule) [Gong et al., 2021], and 4) gastrointestinal polyp segmentation (PolyP) [Jha et al., 2019]. We trained all models using the standard U-Net architecture from MONAI [Cardoso et al., 2022, Kerfoot et al., 2018]. We focused on segmented object size (area) as the downstream clinical metric of interest. Object size is among the most widely adopted quantitative biomarkers across diverse clinical applications [Smith et al., 2003, O’Connor et al., 2008]. For COMPASS-J, which requires a differentiable metric function, we compute area by applying a soft thresholding (a sigmoid function) to the output logits and then summing the resulting probability map. We repeated each experiment over 100 randomized splits to compute average coverage and interval sizes. We provide details on architecture, preprocessing, training, calibration, and testing details in Appendix E. We considered three types of calibration strategies:
1. Output-space calibration: These include split conformal prediction (SCP) [Lei et al., 2018], Conformalized Quantile Regression (Output-CQR) [Romano et al., 2019], and locally adaptive conformal prediction (Local CP) [Lei et al., 2018, Papadopoulos et al., 2008, 2011].
2. End-to-end calibration: We use Conformalized Quantile Regression (E2E-CQR) [Lambert et al., 2024] on a model trained to directly produce pixel-wise lower and upper bounds using the Tversky loss [Salehi et al., 2017].
3. Feature-space calibration: These include COMPASS-L, which directly calibrates the model’s final pre-activation outputs (logits), and COMPASS-J, which constructs prediction intervals by perturbing latent features along dominant directions. As the strong monotonicity is empirically validated in Figures 2 and 12, all our experiments utilize the practical endpoint algorithm. We omitted the original FCP method [Teng et al., 2022] from our main empirical comparison because we found its core adversarial search procedure for finding non-conformity scores to be computationally intractable for large feature spaces and to fail to reliably converge, preventing the computation of defined scores. We instead offer a comparison against a conceptual oracle benchmark for FCP in Table 3: because FCP prediction set is defined by an $L_{p}$ ball in the latent space [Teng et al., 2022] with bounds computed using Linear Relaxation based Perturbation Analysis (LiRPA) [Xu et al., 2020], we empirically find the minimal radius of the $L_{p}$ ball, which, when propagated through the decoding function using LiRPA, achieves the target coverage for the final metric.
Table 1: COMPASS achieves more efficient interval sizes compared to baseline methods across different target coverages. For 4 datasets and 100 random splits, we compare interval lengths at $\alpha=\{0.15,0.1,0.05\}$. We show the output space, end-to-end, and feature calibration methods in red, blue, and green. The shortest mean interval lengths are bolded. For empirical coverages, see Table 2.
[…]
@@
H&E 0.05 4637±630 4408±432 5121±651 6297±722 5646±358 5542±676
H&E 0.10 3160±336 3139±375 3433±293 4223±558 3879±369 3509±333
H&E 0.15 2320±252 2354±146 2679±199 3175±291 2819±207 2550±196
Skin Lesion 0.05 1657±80 1689±83 2569±195 3797±237 10857±65 3273±229
Skin Lesion 0.10 1179±53 1208±58 1351±75 2433±101 4581±36 1813±127
Skin Lesion 0.15 934±30 956±33 943±47 1865±50 2634±44 1124±77
Nodule 0.05 3257±210 3394±280 4150±265 3981±202 7481±46 4589±431
Nodule 0.10 2444±174 2510±180 2788±154 3311±133 5603±57 3076±200
Nodule 0.15 2016±143 2082±142 2150±164 2877±111 4032±64 2408±154
PolyP 0.05 5489±575 6376±769 9162±804 12394±2577 8163±722 8570±766
PolyP 0.10 4056±293 4397±469 6184±616 5965±1011 4981±675 6237±564
PolyP 0.15 3394±290 3686±361 4528±487 4463±481 3913±326 4504±366
@@ 3.1 Standard Conformal Prediction
First, we qualitatively find that COMPASS finds semantically meaningful directions in the latent space that correspond to monotonic changes in metric (Figure 2) for all applications. Next, we compare interval size across methods and datasets for $\alpha=\{0.15,0.1,0.15\}$ (Table 1). We find that output space methods (SCP, CQR, Local) generally achieve valid coverage (Table 2) but yield longer intervals. End-to-end CQR provides tighter intervals. COMPASS (-J and -L) achieves valid coverage with a significantly reduced interval size compared to E2E-CQR. Moreover, we observe an efficiency benefit of using deeper representations (COMPASS-J) compared to logits (COMPASS-L) for the majority of datasets. COMPASS-J generally achieves the shortest interval lengths overall. We also find that COMPASS maintains similar calibration stability as baseline methods (Figure 7). This highlights the efficiency of COMPASS methods.
Figure 2: Visual verification of monotonicity to justify Endpoint-COMPASS. As latent features are shifted along the COMPASS-J direction $\Delta$, the induced segmentation volumes (red contours) monotonically expand. This is the key justification for using our efficient Endpoint-COMPASS in our experimental setup, as it demonstrates mathematical equivalence with the rigorous Envelope-COMPASS. We show a sample from each dataset with perturbation magnitudes $\beta$ targeted at -20%, -10%, 0% (original prediction), +10%, and +20% change in area ($\delta A$). We provide a plot of all metric responses on the testing datasets in Figure 12 and more visual examples in Appendix G.
@@ 3.2 Weighted Conformal Prediction
To empirically validate the effectiveness of WCP under distribution shift, we design an experiment with a controlled, adversarial label shift for the H&E and Skin Lesion datasets. We adopted a dataset-specific partitioning strategy based on sample availability. For H&E, we trained on a restricted subset to prevent data exhaustion, ensuring sufficient minority samples remained. Conversely, for Skin Lesion, we utilized the full dataset, inducing shift by systematically reallocating the class distributions. For H&E, we induced a shift from “easy” to “hard”. We allocated 40% of the Adenocarcinoma samples to the calibration set and the remaining 60% to the test set. This results in a test distribution dominated by difficult samples, leading to baseline undercoverage. For Skin Lesion, we induced a shift from “hard” to “easy”. We allocated 30% of the Melanocytic Nevi (majority/easy) samples to the calibration set, forcing the calibration set to be composed primarily of diverse, difficult lesions. The remaining 70% of the Nevi were allocated to the test set. This resulted in a calibration set that was significantly more difficult than the test set, leading to baseline overcoverage. See Appendix E for more details.
For output-space calibration methods, we used weights based on the ground truth class labels, which we consider as approximately “oracle” weights. Note that the theoretically perfect oracle weights are defined by the density ratio, which precisely corrects for the change in the full, high-dimensional distribution of features between the calibration and test sets. For end-to-end and feature space calibration, we used the latent features [Lambert et al., 2024, Woodland et al., 2023, Anthony and Kamnitsas, 2023] and Jacobians to train auxiliary classifiers using gradient boosting machines [Ke et al., 2017] to distinguish between calibration and test sets. For end-to-end, we do not use Jacobian weights consistent with prior work [Lambert et al., 2024].
Our results (Figure 3) demonstrate that both the choice of weighting information and the choice of calibration layer are critical for achieving robust coverage under these covariate shifts. Simple strategies, such as class-weighting, do not universally recover target coverage; unweighted methods (grey) fail on the H&E dataset, while class-weighted methods (red) fail on Skin Lesion. Furthermore, not all feature-based methods are robust. COMPASS-L and E2E-CQR with feature weighting still fail to maintain coverage under covariate shift on the H&E dataset. In contrast, the COMPASS-J variants with both feature-weighting and Jacobian-weighting were the only methods to consistently maintain the target coverage across both covariate shifts. Among the methods that proved empirically valid, the COMPASS-J variants were also the most efficient. This suggests that 1) the model’s deep features or their Jacobians provide a richer, more adaptive signal for difficulty than simpler class labels or logits, and 2) this deep-layer signal is essential, as methods relying on shallower layers (like COMPASS-L and E2E-CQR) were not robust even when using the same feature-weighting. Our results are tabulated in Table 5.
@@ 4 Discussion
We introduced COMPASS, a unified framework to generate efficient, metric-based prediction intervals by leveraging the inductive biases of neural networks. Under the assumption of exchangeability, COMPASS creates practical feature-space calibration by perturbing a model’s intermediate representations along low-dimensional subspaces maximally sensitive to the target metric. Across four medical segmentation tasks for area estimation, COMPASS achieves valid coverage while producing significantly tighter intervals than traditional output-space and end-to-end baselines for both standard and weighted CP. We discuss several further points below.
COMPASS has several attractive properties. First, it produces instance-adaptive intervals. For COMPASS-J, the perturbation direction for any new sample is found by projecting its Jacobian onto the principal component and reconstructing the direction. Although COMPASS-L applies a uniform scalar shift, it achieves adaptivity because this shift interacts with the spatially varying logit distribution unique to each input, resulting in instance-specific adjustments to the segmentation boundary. Second, COMPASS performs better than naive output-space calibration. In tasks like segmentation, the final metric is a complex non-linear function of the model’s output (the pixel mask). Simply adding a margin to a final object size prediction is a crude approximation that fails to capture how the object size is actually derived from the underlying segmentation. In contrast, COMPASS exploits the model’s internal structure and spatial understanding to directly manipulate and calibrate the metric. COMPASS has more information and degrees of freedom, which allows it to achieve more efficient interval sizes. Third, COMPASS performs better than methods trained using pixel-level losses such as E2E-CQR. Such approaches optimize a proxy (pixel quantiles) for the downstream metric, that may or may not translate into accurate intervals for the final metric. On the other hand, COMPASS is directly calibrated to the final metric, and is therefore more applicable to practical clinical use cases.
Finally, COMPASS is naturally applicable to various segmentation architectures. We also present the interval lengths and coverage for SegResNet [Myronenko, 2018] in Appendix D. COMPASS produces the most efficient interval lengths on the majority of datasets across $\alpha=\{0.05,0.1,0.15\}$ for SegResNet.
@@
COMPASS’s performance is fundamentally dependent on the quality of the pre-trained model’s representations. Our practical algorithm implementation presupposes that the model has learned an inductive bias where the chosen feature space has a reasonably monotonic and sensitive control over the downstream metric. If the features are poorly aligned with the metric, the resulting perturbation direction may be inefficient or non-monotonic, leading to wide or invalid intervals. If so, we recommend using the full discretized sweep across the perturbation range to find the minimum and maximum values. Furthermore, as mentioned in prior work [Lambert et al., 2024], WCP with feature or jacobian weights, is effective for moderate distribution shifts where the calibration and test distributions have significant overlap in the feature space. However, for large shifts, the estimated weights may be inaccurate, which can compromise the validity of the final interval.
[…]
4.1 Practical Considerations
While COMPASS-L may be applicable to UNet-style architectures and object size, other applications may require different setups. Careful design of the different components of COMPASS, specifically COMPASS-J, will significantly enhance its applicability and performance. We provide a list of practical considerations to keep in mind when using COMPASS.
Optimal Feature Layer.
The choice of the feature layer is critical. We find that the optimal layer can be selected empirically by measuring the explained variance of the first principal component ($L=1$) of the Jacobians. High explained variance in the first principal component indicates that the local gradients of the target metric are globally aligned across the samples. Consequently, there is a high chance that a linear traversal along this principal direction results in a consistent, monotonic shift in the output metric for the entire dataset, as the projection scalar maintains a consistent sign. As shown in Figure 5, layers closer to the logits capture the most variance related to object size, validating our use of COMPASS-J. We recommend finding layers with high explained variance of the Jacobians and verify monotonicity with a perturbation sweep (Algorithm 2).
Number of PCA Components.
The number of eigenvectors $L$ controls the dimensionality of the perturbation subspace. We find that for the segmentation area metric, the first principal component already captures over 90% of the metric’s variance (Figure 5). Adding more components does not significantly improve interval tightness (Figure 9). Therefore, we recommend $L=1$ for this task, although this may be task-dependent.
Spatial Entanglement vs Global Semantic Alignment. We observe that the explained variance of the Jacobians is significantly higher when computed on spatially aggregated features compared to raw flattened features (Figure 10). This results from the translation-variant nature of the flattened space, where sensitivity is tied to specific pixel coordinates that vary with object position, leading to orthogonal gradients across samples. In contrast, summing the spatial dimensions isolates the channel-wise semantic sensitivity, which remains consistent across samples. We recommend summing the spatial dimensions of the Jacobians before applying dimensionality reduction techniques.
Computationally Efficiency. COMPASS maintains high computational efficiency, particularly when the metric response is monotonic. The jacobian computation is inexpensive as gradients are calculated with respect to the scalar metric rather than the high-dimensional output image, leveraging autograd. We further resolve potential intractability in PCA by spatially summing the Jacobians, reducing the feature space to a low-dimensional matrix regardless of spatial resolution. While the jacobian and calibration is fast (Table 8 and Figure 11), we recommend precomputing the jacobians for all data before calibration and testing.
LLM Disclosure. During the preparation of this manuscript, we used LLMs (Google Gemini) to polish the initial draft (grammar and style) and improve readability.